OpenStudy (anonymous):

What should be done to solve the following equation? (6.8)x = 34 Multiply each side by 6.8. Subtract 0.8 from both sides, then divide each side by 6. Divide each side by 34. Divide each side by 6.8.

OpenStudy (anonymous):

To solve for x Divide each side by 6.8.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

you would divde both sides by 6.8 to find the value of x, just like cyter said \

OpenStudy (savvy):

\[1 \times x = 34/6.8\]it is \[6.8 \times x = 34\] if we divide by 6.8 on both sides... \[[6.8 \times x]/6.8 = 34/6.8\] which is \[[6.8/6.8] \times x = 34/6.8\] which is \[1 \times x = 34/6.8\] hence x = 34/6.8 so x = 5